# Asymptotics of Proximity Operator for Squared Loss and Performance Prediction of Nonconvex Sparse Signal Recovery

### v3

@@ Abstract
Proximal splitting-based convex optimization is a promising approach to linear inverse problems because we can use some prior knowledge of the unknown variables explicitly. An understanding of the behavior of the optimization algorithms would be important for the tuning of the parameters and the development of new algorithms. In this paper, we first analyze the asymptotic property of the proximity operator for the squared loss function, which appears in the update equations of some proximal splitting methods for linear inverse problems. Our analysis shows that the output of the proximity operator can be characterized with a scalar random variable in the large system limit. Moreover, we apply the asymptotic result to the prediction of optimization algorithms for compressed sensing. Simulation results demonstrate that the MSE performance of the Douglas-Rachford algorithm can be well predicted in compressed sensing with the $\ell_{1}$ optimization. We also examine the behavior of the prediction for the case with nonconvex smoothly clipped absolute deviation (SCAD) regularization.
@@ I Introduction
The above analyses have focused on the performance of the optimizer, whereas the application of CGMT for the performance prediction of optimization algorithms has been considered in [27]. In [27], we have proposed a prediction method for the error performance of the tentative estimate at each iteration of ADMM. Such performance prediction could be utilized to understand the behavior of the algorithm, tune the parameter, and develop a new algorithm. For optimization algorithms other than ADMM, however, such asymptotic performance prediction has not been discussed for linear inverse problems in the literature.
In this paper, we investigate the asymptotic behavior of algorithms using the proximity operator of the squared loss function. As the target optimization problem, we consider the objective function with the squared loss function and a separable regularizer. Since the proximity operator of the squared loss function is not element-wise, we investigate its asymptotic property via the CGMT framework. Specifically, we show that the output of the proximity operator for the squared loss function can be regarded as a scalar random variable in the large system limit. We then utilize the result to propose a prediction method for the evolution of the asymptotic MSE in the Douglas-Rachford algorithm, which is an example of the algorithms using the proximity operator of the squared loss function. Although the derivation of the proposed method is partly non-rigorous, simulation results show that the empirical performance of the Douglas-Rachford algorithm is close to its prediction in compressed sensing with the $\ell_{1}$ optimization. Moreover, we also investigate the behavior for the case with nonconvex regularization in the simulation, though only the convex optimization is considered in [27].
@@ II-B Optimization-Based Approach
in this paper. We refer to $L(\bm{s})=\frac{1}{2}\left\lVert\bm{y}-\bm{A}\bm{s}\right\rVert_{2}^{2}$ as the squared loss function. The function $f_{\lambda}(\cdot):\mathbb{R}^{N}\to\mathbb{R}\cup\left\{+\infty\right\}$ is a regularizer that can incorporate prior knowledge of the unknown vector $\bm{x}$. The parameter $\lambda$ ($>0$) determines the balance between two terms in the objective function.
In the sparse vector reconstruction, for example, the $\ell_{1}$ regularization by $f_{\lambda}(\bm{s})=\lambda\left\lVert\bm{s}\right\rVert_{1}$ is widely used as a convex regularizer. Other than the $\ell_{1}$ regularization, various nonconvex regularizers have been proposed for the sparse vector reconstruction [32]. For example, the smoothly clipped absolute deviation (SCAD) [33] regularizer is given by $f_{\lambda}(\bm{s})=\sum_{n=1}^{N}\tilde{f}_{\lambda}(s_{n})$, where
@@
and $a$ ($>1$) is the parameter.
Various optimization algoirhms have been proposed for the optimization problem in (2). For example, the Douglas-Rachford algorithm [14, 15, 16] solves the optimization problem in (2) by using the proximity operators of $L(\cdot)$ and $f_{\lambda}(\cdot)$. The update equations of the algorithm with the iteration index $k$ ($=0,1,2,\dotsc$) can be written as
@@
where $r_{n}$ and $\left[\mathrm{prox}_{\gamma f_{\lambda}}(\bm{r})\right]_{n}$ are the $n$-th element of $\bm{r}$ and $\mathrm{prox}_{\gamma f_{\lambda}}(\bm{r})$, respectively. As for the SCAD regularizer in (3), the proximity operator is given by
@@
for $a>1+\gamma$. By computing (4) and (5) iteratively, we can obtain a sequence $\left\{\bm{s}^{(k)}\right\}_{k=1,2,\dotsc}$ converging to the solution of the optimization problem in (2) when the regularizer is convex. When the regularizer is nonconvex, the convergence of the sequence is not necessarily guaranteed.
@@ III-A Asymptotics of Proximity Operator for Squared Loss
We assume that each element of the measurement matrix $\bm{A}$ follows a Gaussian distribution in Assumption III.1. This is because we require the Gaussian assumption in the analysis with CGMT [23]. Nonetheless, the universality of random matrices discussed in [34, 35, 36] implies that the analytical result still holds for some other distributions.
@@
has a unique optimizer $(\alpha^{\ast},\beta^{\ast})$. In (11),
@@
By definition in (10), $\hat{\bm{s}}$ is characterized as the solution of the optimization problem in (6). Similarly to [26], we can prove the statement of the proposition by using the standard method with CGMT in [23]. Hence, we do not include the details of the rigorous proof. For the overview of the derivation of (11), see Appendix A. ∎
@@ III-B Application to Performance Prediction of Optimization Algorithm
By using Proposition III.1, we can derive an error prediction method for the estimate $\bm{s}^{(k)}$ in some optimization algorithm. To obtain the prediction method, we make an additional assumption that the regularizer $f_{\lambda}(\cdot)$ is separable as follows.
@@
The function $f_{\lambda}(\cdot):\mathbb{R}^{N}\to\mathbb{R}\cup\left\{+\infty\right\}$ is separable and can be written as $f_{\lambda}(\bm{s})=\sum_{n=1}^{N}\tilde{f}_{\lambda}(s_{n})$ with a function $\tilde{f}_{\lambda}(\cdot):\mathbb{R}\to\mathbb{R}\cup\left\{+\infty\right\}$. We sometimes use the notation $f_{\lambda}(\cdot)$ for the corresponding function $\tilde{f}_{\lambda}(\cdot)$ with the slight abuse of notation.
@@
As an example, we derive the conjecture below for Douglas-Rachford algorithm in (4) and (5). Note that the derivation of the conjecture is partly non-rigorous as [27, Claim III.1].
@@
As discussed in [15], the Douglas-Rachford algorithm is closely related to ADMM analyzed in [27]. However, the relaxation parameters $\rho_{k}$ are included in the Douglas-Rachford algorithm considered in this paper, whereas the corresponding parameters are not considered in [27]. These parameters provide additional flexibility in tuning the algorithm for faster convergence. Furthermore, since we have expressed Proposition III.1 as the result for the proximity operator for the squared loss function, our results have the potential to be extended to other optimization algorithms that utilize the proximity operator.
@@ IV Simulation Results
Here, $\delta_{0}(\cdot)$ is the Dirac delta function, $p_{H}(\cdot)$ is the probability density function corresponding to $\mathcal{N}(0,1)$, and $p_{0}\in(0,1)$ represents the probability of 00. The unknown vector $\bm{x}$ is sparse when $p_{0}$ is large. For the simulations, we assume that the measurement matrix $\bm{A}$ and the noise vector $\bm{v}$ satisfy the conditions in Assumption III.1. For the optimization of $\alpha$ and $\beta$ in (11) and (17), the golden section search is used.
@@ IV-B Distribution of Output of Proximity Operator for Squared Loss
IV-C Performance Prediction of Douglas-Rachford Algorithm with $\ell_{1}$ Regularization
Next, we investigate the validity of Conjecture III.1. For the reconstruction of the sparse vector following (19), we consider the $\ell_{1}$ optimization with $f_{\lambda}(\bm{s})=\lambda\left\lVert\bm{s}\right\rVert_{1}$, i.e.,
@@ IV-C Performance Prediction of Douglas-Rachford Algorithm with $\ell_{1}$ Regularization
The parameters of the Douglas-Rachford algorithm are set as $\gamma=10$ and $\rho_{k}=1$. The initial value of $\bm{z}^{(k)}$ is given by $\bm{z}^{(0)}=\bm{0}$. In Fig. 3, ‘empirical’ refers to the empirical MSE performance of the Douglas-Rachford algorithm in (4) and (5). The MSE curve is the average of the performance for $500$ independent realizations of the reconstruction problem. Also, ‘prediction’ represents the predicted MSE obtained by Conjecture III.1. To calculate the prediction, we compute $(\alpha_{k}^{\ast},\beta_{k}^{\ast})$ and $100,000$ realizations of $(S_{k},Z_{k})$ for $k=0,1,\dotsc$ from (15) and (16), following the similar way to [27]. Here, the expectation in (17) is approximated by using the realizations of $(S_{k},Z_{k})$. In Fig. 3, we also show the asymptotic MSE of the optimizer of (20) as ‘asymptotic MSE of optimizer,’ which is derived by using the standard CGMT approach [23]. The value of the parameter $\lambda$ in (20) is selected to minimize the asymptotic MSE of the optimizer. Figure 3 shows that the empirical performance is close to the prediction when $N$ is sufficiently large. We can also see that they approach the asymptotic MSE of the optimizer after sufficient iterations. To be precise, however, we can observe a slight difference between the prediction and the empirical performance. This is partly because we evaluate the empirical performance for finite $N$, while we assume $M,N\to\infty$ to obtain the asymptotic prediction. Another possible reason is that we approximate $(\alpha_{k}^{\ast},\beta_{k}^{\ast})$ by using many realizations of $(S_{k},Z_{k})$ in the prediction.
Figure 4 shows the MSE performance of the Douglas-Rachford algorithm for different values of $\rho_{k}$.
Figure 4: MSE performance of the Douglas-Rachford algorithm for $\ell_{1}$ optimization ($N=1000$, $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=5.0$).
@@
We then show the MSE performance of the Douglas-Rachford algorithm for different parameters $\gamma$ in Fig. 5.

### v4

@@ Abstract
Proximal splitting-based convex optimization is a promising approach to linear inverse problems because we can use some prior knowledge of the unknown variables explicitly. An understanding of the behavior of the optimization algorithms would be important for the tuning of the parameters and the development of new algorithms. In this paper, we first analyze the asymptotic property of the proximity operator for the squared loss function, which appears in the update equations of some proximal splitting methods for linear inverse problems. Our analysis shows that the output of the proximity operator can be characterized with a scalar random variable in the large system limit. Moreover, we apply the asymptotic result to the prediction of optimization algorithms for compressed sensing. Simulation results demonstrate that the MSE performance of the Douglas-Rachford algorithm can be well predicted in compressed sensing with the $\ell_{1}$ optimization. We also examine the behavior of the prediction for the case with nonconvex smoothly clipped absolute deviation (SCAD) and minimax concave penalty (MCP) regularization.
@@ I Introduction
The above analyses have focused on the performance of the optimizer, whereas the application of CGMT for the performance prediction of optimization algorithms has been considered in [27]. In [27], we have proposed a prediction method for the error performance of the tentative estimate at each iteration of ADMM. Such performance prediction could be utilized to understand the behavior of the algorithm, tune the parameter, and develop a new algorithm. For optimization algorithms other than ADMM, however, such asymptotic performance prediction has not been discussed for linear inverse problems in the literature. Expanding the prediction method to other optimization algorithms would be important for the development of new algorithms and the understanding of the behavior of the algorithms.
In this paper, we investigate the asymptotic behavior of algorithms using the proximity operator of the squared loss function. As the target optimization problem, we consider the objective function with the squared loss function and a separable regularizer. Since the proximity operator of the squared loss function is not element-wise, we investigate its asymptotic property via the CGMT framework. Our main finding is that the optimization problem in the definition of the proximity operator can be analyzed by using CGMT for the squared loss function. Specifically, we show that the output of the proximity operator for the squared loss function can be regarded as a scalar random variable in the large system limit.
By using the analytical result, we also propose a prediction method for the evolution of the asymptotic MSE in proximal splitting methods composed of the proximity operator of the squared loss function and a separable regularizer. As an representative example of such algorithms, we here consider the Douglas-Rachford algorithm in this paper. While the derivation of the proposed method is partly non-rigorous, their simulation results for the $\ell_{1}$ optimization demonstrate their potential as a foundation for more rigorous discussion in future work. These findings would contribute to the broader understanding of optimization algorithms in high-dimensional settings. Moreover, we also investigate the behavior for the case with nonconvex regularization in the simulation, though only the convex optimization is considered in [27]. The effectiveness of the performance prediction across various scenarios highlights its potential to extend to other optimization algorithms, including those involving nonconvex cases.
@@ II-B Optimization-Based Approach
in this paper. We refer to $L(\bm{s})=\frac{1}{2}\left\lVert\bm{y}-\bm{A}\bm{s}\right\rVert_{2}^{2}$ as the squared loss function. The function $f_{\lambda}(\cdot):\mathbb{R}^{N}\to\mathbb{R}\cup\left\{+\infty\right\}$ is a regularizer that can incorporate prior knowledge of the unknown vector $\bm{x}$. The parameter $\lambda$ determines the balance between two terms in the objective function.
In the sparse vector reconstruction, for example, the $\ell_{1}$ regularization by $f_{\lambda}(\bm{s})=\lambda\left\lVert\bm{s}\right\rVert_{1}$ ($\lambda>0$) is widely used as a convex regularizer. The elastic net [32] given by $f_{\lambda}(\bm{s})=\lambda_{1}\left\lVert\bm{s}\right\rVert_{1}+\frac{\lambda_{2}}{2}\left\lVert\bm{s}\right\rVert_{2}^{2}$, where $\lambda_{1},\lambda_{2}$ ($>0$) are the parameters, is also a popular convex regularizer. Other than these convex regularizers, various nonconvex regularizers have been proposed for the sparse vector reconstruction [33]. For example, the smoothly clipped absolute deviation (SCAD) [34] regularizer is given by $f_{\lambda}(\bm{s})=\sum_{n=1}^{N}\tilde{f}_{\lambda}(s_{n})$, where
@@
and $a$ ($>1$) is the parameter. For the minimax concave penalty (MCP) [35] regularizer, the function $\tilde{f}_{\lambda}(s)$ is given by
$\tilde{f}_{\lambda}(s)$ =$=\begin{cases}\lambda\left\lvert s\right\rvert-\dfrac{s^{2}}{2b}&(\left\lvert s\right\rvert\leq b\lambda)\\[5.0pt] \dfrac{b\lambda^{2}}{2}&(\left\lvert s\right\rvert>b\lambda)\end{cases},$ (4)
where $b$ ($>0$) is the parameter.
Various optimization algorithms have been proposed for the optimization problem in (2). For example, the Douglas-Rachford algorithm [14, 15, 16] solves the optimization problem in (2) by using the proximity operators of $L(\cdot)$ and $f_{\lambda}(\cdot)$. The update equations of the algorithm with the iteration index $k$ ($=0,1,2,\dotsc$) can be written as
@@
where $r_{n}$ and $\left[\mathrm{prox}_{\gamma f_{\lambda}}(\bm{r})\right]_{n}$ are the $n$-th element of $\bm{r}$ and $\mathrm{prox}_{\gamma f_{\lambda}}(\bm{r})$, respectively. For the elastic net regularizer $f_{\lambda}(\bm{s})=\lambda_{1}\left\lVert\bm{s}\right\rVert_{1}+\frac{\lambda_{2}}{2}\left\lVert\bm{s}\right\rVert_{2}^{2}$, the proximity operator can be computed as
$\left[\mathrm{prox}_{\gamma f_{\lambda}}(\bm{r})\right]_{n}$ $=\dfrac{\mathrm{sign}(r_{n})\max(\left\lvert r_{n}\right\rvert-\gamma\lambda_{1},0)}{1+\gamma\lambda_{2}}.$ (10)
As for the SCAD regularizer in (3), the proximity operator is given by
@@
for $a>1+\gamma$. Similarly, the proximity operator of the MCP regularizer in (4) can be computed as
$\left[\mathrm{prox}_{\gamma f_{\lambda}}(\bm{r})\right]_{n}$
=$=\begin{cases}0&(\left\lvert r_{n}\right\rvert\leq\gamma\lambda)\\[5.0pt] \dfrac{b}{b-\gamma}(r_{n}-\mathrm{sign}(r_{n})\gamma\lambda)&(\gamma\lambda<\left\lvert r_{n}\right\rvert\leq b\lambda)\\[5.0pt] r_{n}&(\left\lvert r_{n}\right\rvert>b\lambda)\end{cases}$ (12)
for $b>\gamma$. By computing (5) and (6) iteratively, we can obtain a sequence $\left\{\bm{s}^{(k)}\right\}_{k=1,2,\dotsc}$ converging to the solution of the optimization problem in (2) when the regularizer is convex. When the regularizer is nonconvex, the convergence of the sequence is not necessarily guaranteed.
@@ III-A Asymptotics of Proximity Operator for Squared Loss
We assume that each element of the measurement matrix $\bm{A}$ follows a Gaussian distribution in Assumption III.1. This is because we require the Gaussian assumption in the analysis with CGMT [23]. Nonetheless, the universality of random matrices discussed in [36, 37, 38] implies that the analytical result still holds for some other distributions. In [23, Section VIII-F], for example, it is expected that the analysis via CGMT holds for i.i.d. sub-Gaussian matrix. Computer simulations in [23, 27] show that the empirical performance is well predicted even when the measurement matrix is composed of i.i.d. Bernoulli distribution with $p=0.5$.
@@
has a unique optimizer $(\alpha^{\ast},\beta^{\ast})$ [^1] The uniqueness of the solution can be established under certain conditions (e.g., the boundedness of the set of minimizers). However, in the general case, completely removing this assumption proves to be challenging. For further details, please refer to [23, Remark 19].. In (14),
@@
By definition in (13), $\hat{\bm{s}}$ is characterized as the solution of the optimization problem in (7). Similarly to the discussion in [26, Remark IV.1], we can prove (17) by using the standard method with CGMT [23]. Moreover, the second statement of Proposition III.1 is proven in the procedure of [26, Theorem III.2]. Hence, we do not include the details of the rigorous proof. For the overview of the derivation of (14), see Appendix A. ∎
@@ III-B Application to Performance Prediction of Optimization Algorithm
Although the proximity operator itself is not a valid reconstruction method in general, we believe that the result in Proposition III.1 is not only of theoretical interest but also serves as a foundation for understanding and designing optimization algorithms in linear inverse problems. By using Proposition III.1, for example, we can derive an error prediction method for the estimate $\bm{s}^{(k)}$ in some optimization algorithm. To obtain the prediction method, we make an additional assumption that the regularizer $f_{\lambda}(\cdot)$ is convex and separable as follows.
@@
The function $f_{\lambda}(\cdot):\mathbb{R}^{N}\to\mathbb{R}\cup\left\{+\infty\right\}$ is separable and can be written as $f_{\lambda}(\bm{s})=\sum_{n=1}^{N}\tilde{f}_{\lambda}(s_{n})$ with a convex function $\tilde{f}_{\lambda}(\cdot):\mathbb{R}\to\mathbb{R}\cup\left\{+\infty\right\}$. We sometimes use the notation $f_{\lambda}(\cdot)$ for the corresponding function $\tilde{f}_{\lambda}(\cdot)$ with the slight abuse of notation.
@@
To predict the performance of an algorithm by using Proposition III.1, it is necessary that the update equations except for $\mathrm{prox}_{\gamma L}(\cdot)$ are element-wise. As a representative example of such algorithms, we derive the conjecture below for Douglas-Rachford algorithm in (5) and (6). Note that the derivation of the conjecture is partly non-rigorous as [27, Claim III.1].
@@
As discussed in [15], the Douglas-Rachford algorithm is closely related to ADMM analyzed in [27]. In fact, the update equations of ADMM can be regarded as the Douglas-Rachford algorithm for the dual problem of the original optimization problem. However, the relaxation parameters $\rho_{k}$ are included in the Douglas-Rachford algorithm considered in this paper, whereas the corresponding parameters are not considered in the analysis of [27]. These parameters provide additional flexibility in tuning the algorithm for faster convergence. Furthermore, since we have expressed Proposition III.1 as the result for the proximity operator for the squared loss function, our results have the potential to be extended to other optimization algorithms that utilize the proximity operator.
@@ IV Simulation Results
Here, $\delta_{0}(\cdot)$ is the Dirac delta function, $p_{H}(\cdot)$ is the probability density function corresponding to $\mathcal{N}(0,1)$, and $p_{0}\in(0,1)$ represents the probability of 00. The unknown vector $\bm{x}$ is sparse when $p_{0}$ is large. For the simulations, we assume that the measurement matrix $\bm{A}$ and the noise vector $\bm{v}$ satisfy the conditions in Assumption III.1 unless otherwise stated. For the optimization of $\alpha$ and $\beta$ in (14) and (20), the golden section search is used.
@@ IV-B Distribution of Output of Proximity Operator for Squared Loss
IV-C Performance Prediction of Douglas-Rachford Algorithm with Convex Regularization
Next, we investigate the validity of Conjecture III.1. For the reconstruction of the sparse vector following (22), we first consider the $\ell_{1}$ optimization with $f_{\lambda}(\bm{s})=\lambda\left\lVert\bm{s}\right\rVert_{1}$, i.e.,
@@ IV-C Performance Prediction of Douglas-Rachford Algorithm with Convex Regularization
The parameters of the Douglas-Rachford algorithm are set as $\gamma=10$ and $\rho_{k}=1$. The initial values of $\bm{s}^{(k)}$ and $\bm{z}^{(k)}$ are given by $\bm{s}^{(0)}=\bm{z}^{(0)}=\bm{0}$. In Fig. 3, ‘empirical’ refers to the empirical MSE performance of the Douglas-Rachford algorithm in (5) and (6). The MSE curve is the average of the performance for $500$ independent realizations of the reconstruction problem. Also, ‘prediction’ represents the predicted MSE obtained by Conjecture III.1. To calculate the prediction, we compute $(\alpha_{k}^{\ast},\beta_{k}^{\ast})$ and $100,000$ realizations of $(S_{k},Z_{k})$ for $k=0,1,\dotsc$ from (18) and (19), following the similar way to [27]. Here, the expectation in (20) is approximated by using the realizations of $(S_{k},Z_{k})$. In Fig. 3, we also show the asymptotic MSE of the optimizer of (23) as ‘asymptotic MSE of optimizer,’ which is derived by using the standard CGMT approach [23]. The value of the parameter $\lambda$ in (23) is selected to minimize the asymptotic MSE of the optimizer. Figure 3 shows that the empirical performance is close to the prediction when $N$ is sufficiently large. We can also see that they approach the asymptotic MSE of the optimizer after sufficient iterations. To be precise, however, we can observe a slight difference between the prediction and the empirical performance. This is partly because we evaluate the empirical performance for finite $N$, while we assume $M,N\to\infty$ to obtain the asymptotic prediction. Another possible reason is that we approximate $(\alpha_{k}^{\ast},\beta_{k}^{\ast})$ by using many realizations of $(S_{k},Z_{k})$ in the prediction.
We also investigate the performance of the Douglas-Rachford algorithm with the elastic net regularization in Fig. 4.
Figure 4: MSE performance of Douglas-Rachford algorithm with elastic net regularization ($\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.0001$, $\lambda_{1}=\lambda_{2}=0.01$, $\gamma=1$, $\rho_{k}=1$).
In the simulation, we set $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.0001$, $\lambda_{1}=\lambda_{2}=0.01$, $\gamma=10$, and $\rho_{k}=1$. As in the case with the $\ell_{1}$ regularization, the empirical performance is well predicted by the prediction when $N$ is sufficiently large.
Fig. 5 shows the MSE performance of the Douglas-Rachford algorithm with the $\ell_{1}$ regularization for the Bernoulli measurement matrix.
Figure 5: MSE performance of Douglas-Rachford algorithm for $\ell_{1}$ optimization (Bernoulli measurement matrix, $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=10$, $\rho_{k}=1$).
In the evaluation of the empirical performance, the element of the measurement matrix $\bm{A}\in\mathbb{R}^{M\times N}$ is generated from the i.i.d. distribution given by $\Pr(a_{m,n}=1/\sqrt{N})=\Pr(a_{m,n}=-1/\sqrt{N})=0.5$, where $a_{m,n}$ denotes the $(m,n)$ element of $\bm{A}$. For other parameters, we set the same values as in Fig. 3. From Fig. 5, we can see that the empirical performance is well predicted by the prediction even for the Bernoulli measurement matrix, though the Gaussian measurement matrix is assumed in derivation of the prediction.
Figure 6 shows the MSE performance of the Douglas-Rachford algorithm with the $\ell_{1}$ regularization for different values of $\rho_{k}$.
Figure 6: MSE performance of Douglas-Rachford algorithm for $\ell_{1}$ optimization ($N=1000$, $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=5.0$).
@@
We then show the MSE performance of the Douglas-Rachford algorithm with the $\ell_{1}$ regularization for different parameters $\gamma$ in Fig. 7.
@@
We also investigate the performance of the Douglas-Rachford algorithm with the $\ell_{1}$ regularization for different values of $\sigma_{\text{v}}^{2}$ in Fig. 8.
Figure 8: MSE performance of Douglas-Rachford algorithm for $\ell_{1}$ optimization versus $\hat{\sigma}_{\text{v}}^{2}$ ($N=500,\Delta=0.7,p_{0}=0.9,\gamma=10,\rho_{k}=1$).
In the figure, we set $N=500$, $\Delta=0.7$, $p_{0}=0.9$, $\gamma=10$, and $\rho_{k}=1$. From Fig. 8, we can see that the empirical performance is well predicted for different values of $\sigma_{\text{v}}^{2}$. The figure also implies that the performance is improved as the value of $\sigma_{\text{v}}^{2}$ decreases, though the required number of iterations for the convergence increases.